OpenStudy (anonymous):

Im trying to find the derivative. z^(1/2)*e^(-z) Its in the chain rule section Answer is e^(-z)/2z^(1/2)-ze^(1/2-z)

OpenStudy (anonymous):

u gotta use product rule as well

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

\[\frac{d\sqrt{z}e^{-z}}{dz}=\frac{d\sqrt{z}}{dz}e^{-z}+\frac{de^{-z}}{dz}\sqrt{z}\] \[\frac{de^{-z}}{dz}=\frac{de^{-z}}{d(-z)}\frac{d(-z)}{dz} \]